# Can AI Be as Creative as Humans?

### 03 Jan 2024 v1

@@ Abstract
Creativity serves as a cornerstone for societal progress and innovation, but its assessment remains a complex and often subjective endeavor. With the rise of advanced generative AI models capable of tasks once reserved for human creativity, the study of AI’s creative potential becomes imperative for its responsible development and application. This paper addresses the complexities in defining and evaluating creativity by introducing a new concept called Relative Creativity. Instead of trying to define creativity universally, we shift the focus to whether AI can match the creative abilities of a hypothetical human. This perspective draws inspiration from the Turing Test, expanding upon it to address the challenges and subjectivities inherent in evaluating creativity. This methodological shift facilitates a statistically quantifiable evaluation of AI’s creativity, which we term Statistical Creativity. This approach allows for direct comparisons of AI’s creative abilities with those of specific human groups. Building on this foundation, we discuss the application of statistical creativity in contemporary prompt-conditioned autoregressive models. In addition to defining and analyzing a measure of creativity, we introduce an actionable training guideline, effectively bridging the gap between theoretical quantification of creativity and practical model training. Through these multifaceted contributions, the paper establishes a cohesive, continuously evolving, and transformative framework for assessing and fostering statistical creativity in AI models.
@@ 1 Introduction
In this work, we aim to establish a concrete framework for exploring creativity in artificial intelligence. Diverging from traditional methods that attempt to answer the question, “Can AI be creative?”—a question mired in the ambiguous and elusive task of defining creativity absolutely—we instead pivot towards a more concrete inquiry: “Can AI be as creative as humans?” To formally structure this question, we introduce the concept of Relative Creativity (as defined in Definition 1), where an AI model is deemed as creative as a hypothetical, yet realistic, human creator if it can produce works indistinguishable from that creator, as determined by an evaluator. Building on this, we propose a method to evaluate whether and to what degree a model achieves relative creativity through Statistical Creativity (outlined in Theorem 1). This method involves comparing AI outputs with those of actual human creators, coupling with a theoretical guarantee to affirm that statistical creativity is a practical and valid measure for creativity. In line with recent advancements in large language model (LLM), such as GPT-4 (OpenAI,, 2023) and Llama (Touvron et al.,, 2023), our research investigates statistical creativity in auto-regressive models that excel in next-token prediction and generation within a prompting paradigm, as detailed in Corollaries  1 and  2. Furthermore, to connect the theoretical quantification of creativity with practical methods for attaining creativity in model training, we introduce the Statistical Creative Loss (outlined in Corollary 3). This loss function offers a viable strategy for nurturing statistical creativity from training. The corresponding analysis also clarifies the relationship between the log-likelihood of next-token prediction and the necessary volume of training samples required for developing models capable of exhibiting statistical creativity. Overall, our framework presents a comprehensive, evolving methodology for evaluating and enhancing statistical creativity in AI, effectively bridging theoretical concepts with practical application.
@@ 2 Result Overview
In an effort to establish a concrete framework for exploring creativity in artificial intelligence, Section 3 introduces the concept of Relative Creativity (see Definition 1). This concept, echoing the Turing Test’s comparative approach to intelligence assessment, proposes that an AI model can be deemed as creative as a hypothetical human creator if its outputs are indistinguishable from those of the creator, as judged by a designated evaluator. Applying this concept directly to AI model evaluation encounters practical challenges, mainly due to the hypothetical nature of the creator, which hinders empirical validation. To overcome this, Section 4.1 presents Statistical Creativity (Theorem 1), a methodology that makes assessing Relative Creativity both feasible and realistic, anchoring it in the context of existing individuals’ creative outputs. Further exploration in Section 4.2 delves into autoregressive models, a prevalent technique in generative models, introducing Autoregressive Statistical Creativity (Theorem 2). Moreover, Section 4.3 applies this framework to cutting-edge model paradigm, like prompt-conditioned large language models (LLMs), as explored in Brown et al., (2020). This section unveils Prompt-Contextualized Autoregressive Statistical Creativity (Corollary 2). In Section 5, we introduce the Statistical Creative Loss (Corollary 3), bridging the theoretical-practical gap and providing actionable guidelines for training creative AI models. The analysis here clarifies the relationship between the upper bound of next-token prediction log-likelihood and the required volume of training samples for achieving Statistical Creativity. Our findings highlight that the quantity of creator-creation pairs are crucial, rather than merely having a large volume of creation data. This insight renders the concept of Statistical Creativity particularly applicable in current AI frameworks based on next-token prediction (as highlighted in Remark 4). We strive for this paper not only to offer theoretical perspectives but to guide the discourse on AI creativity, advocating for relative evaluations to spur empirical research and establish a framework for assessing and augmenting the creative capabilities of AI models.
We summarize our contributions as follows.
@@
Introduction of Relative Creativity: We present the concept of Relative Creativity, a notion of evaluating the creative capabilities of artificial intelligence (AI) systems. This concept focuses on a comparative evaluation, eschewing absolute definitions in favor of a more relative understanding of creativity. Integration of Subjectivity in Creative Assessment: Relative creativity acknowledges the inherent subjectivity in creativity and incorporates the subjectivity into the comparison process, akin to how the Turing Test assesses intelligence through comparison rather than fixed definitions. The subjective nature of creativity is crystallized in the choice of the anchor population. This allows the study of AI creativity to maintain a degree of objectivity. Conceptualization of Statistical Creativity: To facilitate empirical studies, we introduce Statistical Creativity, a means for evaluating whether and to what extent an AI model achieves relative creativity by assessing its ability to replicate the creative abilities of a specific human population. Application to Autoregressive Models: Delving into autoregressive models, we propose a practical measure of statistical creativity for autoregressive models with next-token prediction. This measure can be applied to prompt-conditioned autoregressive models, ensuring its applicability to cutting-edge Large Language Models (LLMs). Development of Statistical Creativity Loss: We propose the Statistical Creativity Loss, which aligns theoretical principles with practical training guidelines for fostering creative AI, and provides insights into collecting creator-creation data and optimizing training objectives.
@@ 3.1 What is Relative Creativity?
Diverging from traditional methods setting an absolute threshold or checklist to determine if AI can be deemed creative, that necessitates a foray into the contentious task of defining creativity universally, we shift the study by proposing an creativity notion for AI called Relative Creativity. Relative creativity sidesteps the difficulties associated with defining creativity in absolute terms, reminiscent the Turing Test’s approach to intelligence evaluation (Turing,, 2009). Note, the Turing Test eschews absolute definitions of intelligence, instead opting for a relative metric that contrasts machine behavior with human responses in conversational scenarios. In the evaluation of creativity, an AI model is deemed “relatively creative” if it can generate creations that are indistinguishable from those of a hypothetical yet plausible human creator, as judged by an evaluator. In simpler terms, relative creativity posits that an AI model is as creative as a hypothetical human if its creations, when conditioned on the biographical data or characteristics of this hypothetical creator, are indistinguishable from what that human would have created. The way in which this notion of creativity is “relative” is that it depends on the individual to which the entity is being compared. For instance, an AI system might seem highly creative when compared to a non-expert human creator but may seem less so when compared to a expert designer or artist. Relative creativity distinguishes itself through acknowledges the inherently subjective facets of creativity—such as originality, divergent thinking, and problem-solving skills—these elements are integrated into the anchor selection process. Because the subjectivity of creativity is crystallized into the choice of the anchor human, relative creativity leaves the study of AI creativity to be objective. Regardless of whether the creative outputs of a specific individual or group satisfy the varied, subjective standards for what constitutes creativity, this notion reframes the debate, enabling a focused, empirical investigation into whether an AI model can replicate the creative capacities of a predefined human benchmark.
@@ 3.2 Notation and Formal Definition
While we can theoretically ascertain the value of $L(\mathcal{M}(\bar{c}),\bar{c})$, its practical assessment becomes intricate due to the inaccessibility of $\bar{c}$ — given that its corresponding human doesn’t exist. To address this, a theoretical framework that can bridge $L(\mathcal{M}(\bar{c}),\bar{c})$ with observable real-world data is required, enabling us to empirically measure whether AI possesses the relative creativity.
4 Measuring Relative Creativity
In this section, we introduce a methodology to quantify AI’s creative potential and establish a measurable framework. This approach evaluates whether and to what extent a model achieves Relative Creativity, basing the assessment on observable outputs of existing human creators.
4.1 Statistical Creativity
Consider the evaluation metric $E_{0}(q)$, an empirical measure designed to estimate the indistinguishability between the creative abilities of an AI model and those of human creators. The human creators are represented as $(c_{i})_{i=1}^{n}\sim(\mathcal{D}_{C})^{\otimes n}$, $c_{i}$ is independent and identically sampled from creator distribution $\mathcal{D}_{C}$. Formally, $E_{0}(q)$ is defined as:
@@ 4.1 Statistical Creativity
To facilitate the adjustment from theoretical creators (as defined in Definition 1) to actual creators, whose data is readily obtainable, we present the concept of Statistical Creativity (outlined in Theorem 1). This theorem explicitly outlines the conditions under which an AI model can be classified as exhibiting $\delta$-creativity, in relation to a specific group of human creators, as assessed by evaluator $L$. Specifically, if $E_{0}(q)<\delta$ and we have a sufficiently large sample set for evaluation $\left(n\geq\frac{\ln(1/\delta^{\prime})}{2(\delta-E_{0})^{2}}\right)$, then the AI can be classified as $\delta$-creativity as the humans from $\mathcal{D}_{C}$.
@@
Moving beyond a “creative or not” categorization, statistical creativity presents a sophisticated perspective where creativity is characterized by its deviation, $\delta$, to a pre-selected existing human creators as benchmarks. Crucially, statistical creativity does not require an AI to perfectly replicate human creators. Instead, it highlights the importance of achieving a certain level of similarity that is appreciable through the lens of the evaluator, represented by $L$. When a human assumes the role of the evaluator, a creatively successful AI is expected to skillfully mimic a novel creator, as perceived from the human evaluator’s perspective.
4.2 Measuring Statistical Creativity of Autoregressive Model
In this section, we delve into applying Statistical Creativity to autoregressive models, which are fundamental to large language models (LLMs) recognized for exhibiting a certain level of creative capabilities (Zhao et al.,, 2023). We specifically explore the specific form of Statistical Creativity in this context, thereby enhancing our understanding of it within next-token prediction mechanisms.
@@ 4.2 Measuring Statistical Creativity of Autoregressive Model
It’s worth noting that by setting $T=1$, the non-auto-regressive structure is retained, preserving the metric’s versatility. In addition, an intriguing aspect of $E_{1}(q)$ is the term $\frac{1}{r(c_{i})}$. This term plays a pivotal role in ensuring that if the entropy in the creation process by a creator $c_{i}$ is vast, the corresponding log-likelihood receives lesser weight. This adjustment effectively captures the inherent unpredictability and diversity characterizing human creativity.
@@
Within this context, $\tau>0$ stands as the threshold fulfilling the condition stipulated in Assumption 1. Furthermore, we introduce $r_{\min}$ as the least positive real number ensuring $r_{\min}\leq r(c)$ with near certainty for $c\sim\mathcal{D}_{C}$. The above assumption delineates the conditions under which an AI is deemed statistically creative. With the condition, the following theorem presents the measurement of statistical creativity in the next-token prediction setup.
@@
Contrasting it with Theorem 1, Theorem 2 reveals that demonstrating creativity does not necessarily require strictly imitating each creator. Instead, it emphasizes the significance of the log-likelihood in next-token prediction across a wide range of creators. Furthermore, Theorem 2 addresses a critical question: Is it essential to have a large number of samples from each creator to demonstrate a model’s statistical creativity? This inquiry is vital, considering the potential bottleneck in creative AI development due to the extensive collection of each creator’s works. Theorem 2 provides insight into this issue, indicating that amassing a vast dataset for each creator is not as crucial as previously thought. What is more important, as Theorem 2 highlights, is the diversity of creator-creation pairs (i.e., $n$ in $\{(x_{i},c_{i})\}_{i=1}^{n}$). This implies that with a sufficiently large number of creators, the sample size per creator can be relatively small, thereby alleviating data collection challenges in developing creative AI models.
4.3 Measuring Statistical Creativity of Large Language Model
The technique of prompting stands as a potent tool for unlocking the capabilities inherent in foundation models, especially Large Language Models (LLMs) (Bommasani et al.,, 2022). In modeling this prompting setup, we denote $U$ as a set encompassing various contexts of creations. Each context within this set is denoted as $u$. We define $\mathcal{D}_{U}$ as the probability distribution over $U$. Define $\mathcal{D}_{U,C}(u,c)=\mathcal{D}_{C}(c)\mathcal{D}_{U}(u)$, where $\{u_{i}\}_{i=1}^{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}(\mathcal{D}_{U})^{\otimes n}$ symbolizes a sequence of existing context prompts, assumed to be independently and identically distributed. The notation $\bar{u}\sim\mathcal{D}_{U}$ corresponds to a possible new context prompt. To clarify the distinction between the variables $c$ and $u$, consider a practical example involving GPT-4 (OpenAI,, 2023). Here, the creator’s information, such as a biography, is represented by $c$ and serves as the system prompt, while $u$ corresponds to the user’s input prompt.

### 12 Jan 2024 v3

@@ Abstract
Creativity serves as a cornerstone for societal progress and innovation. With the rise of advanced generative AI models capable of tasks once reserved for human creativity, the study of AI’s creative potential becomes imperative for its responsible development and application. In this paper, we provide a theoretical answer to the question of whether AI can be creative. We prove in theory that AI can be as creative as humans under the condition that AI can fit the existing data generated by human creators. Therefore, the debate on AI’s creativity is reduced into the question of its ability of fitting a massive amount of data. To arrive at this conclusion, this paper first addresses the complexities in defining creativity by introducing a new concept called Relative Creativity. Instead of trying to define creativity universally, we shift the focus to whether AI can match the creative abilities of a hypothetical human. This perspective draws inspiration from the Turing Test, expanding upon it to address the challenges and subjectivities inherent in assessing creativity. This methodological shift leads to a statistically quantifiable assessment of AI’s creativity, which we term Statistical Creativity. This concept allows for comparisons of AI’s creative abilities with those of specific human groups, and facilitates the theoretical findings of AI’s creative potential. Building on this foundation, we discuss the application of statistical creativity in prompt-conditioned autoregressive models, providing a practical means for evaluating creative abilities of contemporary AI models, such as Large Language Models (LLMs). In addition to defining and analyzing creativity, we introduce an actionable training guideline, effectively bridging the gap between theoretical quantification of creativity and practical model training. Through these multifaceted contributions, the paper establishes a framework for understanding, evaluating and fostering creativity in AI models.
@@ 1 Introduction
In this work, we aim to establish a concrete framework for exploring creativity in artificial intelligence. Diverging from traditional methods that attempt to answer the question, “Can AI be creative?”—a question mired in the ambiguous and elusive task of defining creativity absolutely—we instead pivot towards a more concrete inquiry: “Can AI be as creative as humans?” Theoretically, we demonstrate that this is feasible if AI can properly fit data produced by human creatives. In order to reach this conclusion, we formally structure the problem by introducing the concept of Relative Creativity (as defined in Definition 1), where an AI model is deemed as creative as a hypothetical, yet realistic, human creator if it can produce works indistinguishable from that creator, as determined by an evaluator. Building on this, we present Statistical Creativity (outlined in Theorem 1), a means for understanding whether and to what degree AI achieves creativity by comparing its outputs with those of actual human creators. Through the lens of statistical creativity, we theoretically demonstrate that AI can be as creative as human beings under the condition that AI fits and learns from existing data produced by human creators. This implies that AI can learn from past data and evolve into an entity capable of generating creative content that, while not yet existing, could conceivably be produced by a new human creator in the future. In line with recent advancements in large language models (LLMs), such as GPT-4 (OpenAI,, 2023) and Llama (Touvron et al.,, 2023), our research investigates statistical creativity in autoregressive models that excel in next-token prediction and generation within a prompting paradigm, as detailed in Corollaries  1 and  2. These corollaries provide guidance into the practical and valid measure for creativity of current AI models. Furthermore, to connect the theoretical quantification of creativity with practical methods for attaining creativity in model training, we introduce the Statistical Creative Loss (outlined in Corollary 3). This loss function offers a viable strategy for nurturing statistical creativity during training. The corresponding analysis also clarifies the relationship between the log-likelihood of next-token prediction and the necessary volume of training samples required for developing models capable of exhibiting statistical creativity. Overall, our framework offers a comprehensive and evolving framework for the theoretical exploration, practical evaluation, and enhancement of creativity in AI, effectively bridging theoretical concepts with practical application.
@@ 2 Result Overview
In an effort to establish a concrete framework for exploring creativity in artificial intelligence, Section 3 introduces the concept of Relative Creativity (see Definition 1). This concept, mirroring the Turing Test’s comparative approach to intelligence assessment, suggests that an AI model can be deemed as creative as a hypothetical human creator if its outputs are indistinguishable from those of the creator, as judged by a designated evaluator. To facilitate the theoretical exploration of AI’s creativity, Section 4.1 presents Statistical Creativity (Theorem 1), a method that anchors relative creativity in relation to the outputs of existing individuals. We theoretically demonstrate that AI possesses the potential for creativity comparable to that of humans, provided it can assimilate data produced by human creators. This theoretical result reduces the debate on AI’s creativity to the question of its ability to fit a massive amount of data (Remark 5). Further exploration in Section 4.2 delves into autoregressive models, a prevalent technique in generative AI, introducing Autoregressive Statistical Creativity (Theorem 2). Moreover, Section 4.3 applies this framework to cutting-edge model paradigm, like prompt-conditioned large language models (LLMs), as explored in Brown et al., (2020). This section unveils Prompt-Contextualized Autoregressive Statistical Creativity (Corollary 2). In Section 5, we introduce the Statistical Creative Loss (Corollary 3), bridging the theoretical-practical gap and providing actionable guidelines for training creative AI models. The analysis here clarifies the relationship between the upper bound of next-token prediction log-likelihood and the required volume of training samples for achieving statistical creativity. Our findings highlight that the quantity of creator-creation pairs are crucial, rather than merely having a large volume of creation data. We strive for this paper not only to offer theoretical perspectives but to guide the discourse on AI creativity, advocating for relative evaluations to spur empirical research and establish a framework for assessing and augmenting the creative capabilities of AI models. We summarize our contributions as follows.
@@
Theoretical Result on AI’s Creativity: We theoretically show that AI can achieve a level of creativity comparable to that of humans under the condition that AI can properly fit the existing data created by human beings. This implies that AI, through its ability to learn from historical data, has the potential to evolve into an entity capable of generating novel, creative outputs. These outputs, while not currently existing, are likely to be generated by a new human creator in the future. Introduction of Relative Creativity: We present the concept of Relative Creativity, a notion for understanding the creative capabilities of artificial intelligence (AI). This concept is defined from a comparative perspective, eschewing absolute definitions in favor of a relative understanding of creativity. Integration of Subjectivity: Relative creativity acknowledges the inherent subjectivity in creativity and incorporates the subjectivity into the comparison process, akin to how the Turing Test assesses intelligence through comparison rather than fixed definitions. The subjective nature of creativity is crystallized in the choice of the anchor population, allowing the study of AI creativity to maintain a degree of objectivity. Introduction of Statistical Creativity: To facilitate theoretical findings, we introduce Statistical Creativity, a means for assessing whether and to what extent an AI model achieves creativity by comparing its ability to those of a specific human population. Application in Contemporary AI Models: Delving into autoregressive models, we propose a practical measure of statistical creativity for autoregressive models with next-token prediction. This measure can be applied to prompt-conditioned autoregressive models, ensuring its applicability to cutting-edge Large Language Models (LLMs). Practical Training Guidelines: We propose the Statistical Creativity Loss, which aligns theoretical principles with practical training guidelines for fostering creative AI, and provides insights into collecting creator-creation data.
@@ 3.1 What is Relative Creativity?
Diverging from traditional methods setting an absolute threshold or checklist to determine if AI can be deemed creative, that necessitates a foray into the contentious task of defining creativity universally, we shift the study by proposing an creativity notion for AI called Relative Creativity. Relative creativity sidesteps the difficulties associated with defining creativity in absolute terms, reminiscent the Turing Test’s approach to intelligence evaluation (Turing,, 2009). Note, the Turing Test eschews absolute definitions of intelligence, instead opting for a relative metric that contrasts machine behavior with human responses in conversational scenarios. In the evaluation of creativity, an AI model is deemed “relatively creative” if it can generate creations that are indistinguishable from those of a hypothetical yet plausible human creator, as judged by an evaluator. In simpler terms, relative creativity posits that an AI model is as creative as a hypothetical human if its creations, when conditioned on the biographical data or characteristics of this hypothetical creator, are indistinguishable from what that human would have created. The way in which this notion of creativity is “relative” is that it depends on the individual to whom the entity is being compared. For instance, an AI system might seem highly creative when compared to a non-expert human creator but may seem less so when compared to a expert designer or artist. Relative creativity distinguishes itself through acknowledges the inherently subjective facets of creativity—such as originality, divergent thinking, and problem-solving skills—these elements are integrated into the anchor selection process. Because the subjectivity of creativity is crystallized into the choice of the anchor human, relative creativity leaves the study of AI creativity to be objective. Regardless of whether the creative outputs of a specific individual or group satisfy the varied, subjective standards for what constitutes creativity, this notion reframes the debate, enabling a focused, empirical investigation into whether an AI model can replicate the creative capacities of a predefined human benchmark.
@@ 3.2 Notation and Formal Definition
4 Statistical Creativity
In this section, we introduce the Statistical Creativity to enable the theoretical exploration of AI’s creativity. This approach assesses whether and to what extent a model achieves creativity, basing the comparative assessment on observable outputs of existing human creators.
4.1 Statistical Creativity of AI
Consider the metric $E_{0}(q)$, a measure designed to assess the indistinguishability between the creative abilities of an AI model and those of human creators. The human creators are represented as $(c_{i})_{i=1}^{n}\sim(\mathcal{D}_{C})^{\otimes n}$, $c_{i}$ is independent and identically sampled from creator distribution $\mathcal{D}_{C}$. Formally, $E_{0}(q)$ is defined as:
@@ 4.1 Statistical Creativity of AI
To facilitate the adjustment from theoretical creators (as defined in Definition 1) to actual creators, whose data is readily obtainable, we present the concept of Statistical Creativity (outlined in Theorem 1). This theorem explicitly outlines the conditions under which an AI model can be deemed as exhibiting $\delta$-creativity in relation to a specific group of human creators, as assessed by evaluator $L$. Specifically, if $E_{0}(q)<\delta$ and we have a sufficiently large sample set for evaluation $\left(n\geq\frac{\ln(1/\delta^{\prime})}{2(\delta-E_{0})^{2}}\right)$, then the AI can be regarded as $\delta$-creative as humans from $\mathcal{D}_{C}$.
@@
In this theorem, we theoretically demonstrate that AI can be as ($\delta$-)creative as humans, provided it can assimilate a specified amount of data ($n$) from human creators. Additionally, Moving beyond a “creative or not” categorization, statistical creativity presents a sophisticated perspective where creativity is characterized by its deviation, $\delta$, to a pre-selected existing human creators as benchmarks. Crucially, statistical creativity does not require an AI to perfectly replicate human creators. Instead, it highlights the importance of achieving a certain level of similarity that is appreciable through the lens of the evaluator, represented by $L$. When a human assumes the role of the evaluator, a creatively successful AI is expected to skillfully mimic a novel creator, as perceived from the human evaluator’s perspective.
4.2 Statistical Creativity for Autoregressive Model
In this section, we delve into applying statistical creativity into autoregressive models, which are fundamental to the contemporary Generative AI models. We explore the specific form of statistical creativity in this context, thereby enhancing our understanding of it within next-token prediction mechanisms.
@@ 4.2 Statistical Creativity for Autoregressive Model
It’s worth noting that by setting $T=1$, the non-autoregressive structure is retained, preserving the metric’s versatility. In addition, an intriguing aspect of $E_{1}(q)$ is the term $\frac{1}{r(c_{i})}$. This term plays a pivotal role in ensuring that if the entropy in the creation process by a creator $c_{i}$ is vast, the corresponding log-likelihood receives lesser weight. This adjustment effectively captures the inherent unpredictability and diversity characterizing human creativity.
@@
Within this context, $\tau>0$ stands as the threshold fulfilling the condition stipulated in Assumption 1. Furthermore, we introduce $r_{\min}$ as the least positive real number ensuring $r_{\min}\leq r(c)$ with near certainty for $c\sim\mathcal{D}_{C}$. The above assumption delineates the conditions under which an AI is deemed statistically creative. With the condition, the following theorem presents the statistical creativity in the next-token prediction setup.
@@
Contrasting it with Theorem 1, Theorem 2 reveals that demonstrating creativity does not necessarily require strictly imitating each creator. Instead, it emphasizes the significance of the log-likelihood in next-token prediction across a wide range of creators. Furthermore, Theorem 2 addresses a critical question: Is it essential to have a large number of samples from each creator to demonstrate a model’s statistical creativity? This inquiry is vital, considering the potential bottleneck in creative AI development due to the extensive collection of each creator’s works. Theorem 2 provides insight into this issue, indicating that amassing a vast dataset for each creator is not as crucial as previously thought. What is more important, as Theorem 2 highlights, is the creator-creation pairs (i.e., $n$ in $\{(x_{i},c_{i})\}_{i=1}^{n}$). This implies that with a sufficiently large number of creators, the sample size per creator can be relatively small, thereby alleviating data collection challenges in developing creative AI models.
4.3 Statistical Creativity for Large Language Model
The technique of prompting serves as a powerful tool for harnessing the inherent capabilities of foundation models, particularly Large Language Models (LLMs) (Bommasani et al.,, 2022), which are recognized as exhibiting a certain level of creative potential (Zhao et al.,, 2023). In modeling this prompting setup, we denote $U$ as a set encompassing various contexts of creations. Each context within this set is denoted as $u$. We define $\mathcal{D}_{U}$ as the probability distribution over $U$. Define $\mathcal{D}_{U,C}(u,c)=\mathcal{D}_{C}(c)\mathcal{D}_{U}(u)$, where $\{u_{i}\}_{i=1}^{n}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}(\mathcal{D}_{U})^{\otimes n}$ symbolizes a sequence of existing context prompts, assumed to be independently and identically distributed. The notation $\bar{u}\sim\mathcal{D}_{U}$ corresponds to a possible new context prompt. To clarify the distinction between the variables $c$ and $u$, consider a practical example involving GPT-4 (OpenAI,, 2023). Here, the creator’s information, such as a biography, is represented by $c$ and serves as the system prompt, while $u$ corresponds to the user’s input prompt.